# LastingBench: Defend Benchmarks Against Knowledge Leakage

### v1

@@ Abstract
The increasing complexity of large language models (LLMs) raises concerns about their ability to “cheat” on standard Question Answering (QA) benchmarks by memorizing task-specific data. This undermines the validity of benchmark evaluations, as they no longer reflect genuine model capabilities but instead the effects of data leakage. While prior work has focused on detecting such leakage, little attention has been given to mitigating its impact and preserving the long-term utility of benchmarks. In this paper, we introduce LastingBench, a novel framework designed to continuously reinforce and safeguard existing benchmarks against knowledge leakage. LastingBench identifies leakage points in the context through perturbation, then rewrites the leakage points to counterfactual ones-disrupting memorization while preserving the benchmark’s original evaluative intent. Evaluations of state-of-the-art QA benchmarks show significant performance gaps, highlighting the efficacy of LastingBench in reducing memorization effects. LastingBench offers a practical and scalable solution to ensure benchmark robustness over time, promoting fairer and more interpretable evaluations of LLMs. Our code and data are available at https://github.com/Seriousss/LastingBench.
@@
Yixiong Fangthanks: Equal contribution.   Tianran Sun11footnotemark: 1   Yuling Shi   Min Wang   Xiaodong Guthanks: Correspondence: xiaodong.gu@sjtu.edu.cn Shanghai Jiao Tong UniversityUniversity of Pennsylvania {fangyixiong,seriousss,yuling.shi,xiaodong.gu}@sjtu.edu.cn minwg27@gmail.com
@@ 1 Introduction
The rapid advancement of LLMs has introduced critical challenges to the reliability and validity of QA evaluation benchmarks. Due to the opaque and large-scale nature of LLM training pipelines, these models often memorize parts of benchmark datasets Xu et al. (2024c); Balloccu et al. (2024); Geva et al. (2023); Deng et al. (2023); Cheng et al. (2025). This unintended data leakage enables models to “cheat” during evaluation, producing correct answers without genuine understanding. As a result, evaluation scores on standard benchmarks may no longer reflect genuine model capabilities, but instead represent latent memorization effects. This poses a serious threat to fair and meaningful measurement in NLP research Sainz et al. (2023a); Oren et al. (2023); Shi et al. (2023); Dekoninck et al. (2024b).
@@
In this paper, we propose LastingBench , a novel framework to reinforce and safeguard existing benchmarks against knowledge leakage, specifically targeting the long-context QA domain. Rather than continuously building new benchmarks, LastingBench enhances existing ones by identifying leakage points and rewriting them using counterfactuals, preserving the benchmark’s original purpose while disrupting model memorization. Our methodology begins with perturbation-based detection. By systematically perturbating the original context and the question, we assess whether a more relies on internal memorization rather than contextual clues. When leakage is detected, we locate critical evidence segments—the minimal context required to justify the answer—using enriched chain-of-thought (CoT) queries from an LLM. These segments are then rewritten into plausible yet semantically contradictory counterfactuals. This makes the benchmark robust against memorization while maintaining its reasoning challenge.
@@ 2.1 Study Design
1) Context Perturbation: We exclude the context and prompt the model to answer the original question barely on the question. If the model can still provide the correct answer, i.e., $\mathcal{L}(q)=a^{\star}$. This suggests its reliance on internal knowledge, indicating knowledge leakage.
2) Question Perturbation: We either rephrase the original question $q$ into a semantically equivalent form $\tilde{q}$ or reformulate it into a logically contradictory version $q_{con}$, where the correct answer should conflict with the original. If the model answers the original question correctly but fails on the rephrased version, $\mathcal{L}(\tilde{q},\mathbf{C})\neq a^{\star}$, or provide the original answer even when the question is contradictory, $\mathcal{L}(q_{\text{con}},\mathbf{C})=a^{\star}$, this suggests dependence on memorized internal knowledge rather than contextual comprehension.
@@ 2.2 Experimental Setup
Studied Models We select a diverse set of models with varying parameter sizes and architectures for our experiments.(1) Both open-source models (Qwen2.5-32B-Instruct Qwen (2025), Llama-3.1-8B-Instruct et al (2024), Phi-4 Abdin et al. (2024)) and proprietary SOTA models including GPT-4o Hurst et al. (2024) and DeepSeek-v3 Liu et al. (2024) are evaluated. (2) For the same model family, we also select models with different versions (e.g., Qwen3-32B vs. Qwen2.5-32B-Instruct) and parameter sizes(e.g., Qwen2.5-8B-Instruct vs. Qwen2.5-32B-Instruct), enabling multi-dimension comparisons.
@@ 2.3.2 Results by Question Perturbation
The results shown in Table 2 reveal varying levels of memorization across models. Notably, Claude-3-haiku shows the highest overlap on HotpotQA, with 40% of contradictory questions yielding the same answer as the original, despite the altered semantics. Similar trends are observed for models like Qwen2.5-32B-Instruct and Phi-4. Such findings suggest that current LLMs heavily rely on memorized internal answers rather than reasoning over the given context.
@@ 2.4 Current Landscape and Trend
Our findings reveal that knowledge leakage is already widespread across long-context benchmarks, and the situation is deteriorating. Newer and larger models such as Qwen3 exhibit much higher leakage than their predecessors, as evidenced by drastic performance drops under question rephrasing. This trend suggests that mainstream benchmarks gradually no longer reliably reflect models’ genuine ability—instead, they often measure memorization and result in inflated numbers. These all highlight the urgent need for a pipeline to restore and defend benchmarks.
@@
21:     return Merge($C,\,C_{\text{crit}},\,C^{(i^{\star})}_{\text{cf}}$)
@@ 3.2 Counterfactual Rewriting
To select the most effective conterfactual, we compute the conditional perplexity of the original question $q$ as:
@@
where, $\text{PPL}(q)$ is the perplexity of the question $q$ answered without context, and $\text{PPL}(q\mid\mathbf{C}_{cf}^{(i)})$ measures perplexity given the i-th conterfactual. A higher CPPL indicates a stronger contradiction to the model’s internal knowledge. We select the counterfactual with the highest conditional perplexity:
@@ 4.1 Revised Dataset Details
We utilize DeepSeek-R1 as the leakage detecting model, evaluating with same methods in Section 2. Table 4 reports the percentages of revised (leaked) and unchanged entries for each dataset used in our experiments.
Dataset Revised Entries Unchanged Enrtries
@@ 4.2 Efficacy in Model Evaluation
As is displayed in Table 5, across all datasets, most models exhibit a consistent performance decline across different datasets. The performance drop is more pronounced on HotpotQA and 2WikiMQA, with HotpotQA showing the most significant decline (30% for Deepseek-v3). This aligns with our observations in §2.3.1 regarding context perturbation-based leakage detection. Among all evaluated models, Claude-3-Haiku (27% on HotpotQA) and the Qwen 3 series (30% on 2WikiMQA) show more pronounced performance degradation on the revised datasets, which is consistent with the leakage detection results presented in §2.3.2. This may be attributed to the fact that these models are relatively newer and therefore more likely to have been exposed to these datasets during the training stage, leading to potentially inflated performance. In addition, larger models tend to exhibit a higher degree of data memorization within their parameters. For example, GPT-4o and Deepseek-v3 shows a consistent decline on all the datasets. In contrast, our revised datasets provide a more trustworthy assessment of the models’ genuine long-context reasoning ability.
@@ 4.3 Comparative Analysis
Table 6 shows the model performance on HotpotQA. The results indicate that model performance on the randomly reformulated dataset falls short of the original, yet surpasses that on the counterfactual dataset, which suggests that our counterfactual rewriting method is more resistant to model cheating and poses a greater challenge, making it less likely to be memorized during pretraining. As a result, it offers a more faithful and enduring evaluation of the model’s long-context reasoning abilities.
@@ 4.4 Efficacy in Model Training
To examine whether our dataset defense effectively increase the learning complexity for language models and consequently, reduce the likelihood of models memorizing information during training stage, we conducted an empirical evaluation. Specifically, we fine-tuned four distinct models on both the original 2WikiMQA dataset and our revised dataset [^3] Details of our fine-tuning settings are provided in Section A. Introducing knowledge conflicts in datasets naturally creates greater complexity during the learning process. Ideally, this discourages models from memorizing knowledge.
As is displayed in Figure 4, throughout the finetuning process, all models exhibit higher training loss on our revised dataset than on the original dataset, indicating that it exposes them to knowledge conflict situations. The consistently higher training loss values indicate that the revised data poses greater learning difficulty, which in turn mitigates the risk of future data leakage by making it less likely for models to memorize or retain such information.
@@ 5.1 Contamination Detection
The evaluation of Large Language Models (LLMs) is increasingly complicated by data contamination—the unintended overlap between training data and evaluation benchmarks (Fu et al., 2025a; Cheng et al., 2025; Xu et al., 2024a). This issue can artificially inflate performance scores, misrepresenting LLMs’ true generalization capabilities. Vast pre-training datasets, often web-scraped, frequently include common benchmark samples, leading to deceptively high leaderboard scores that may not reflect real-world robustness (Tonmoy et al., 2024; Xu et al., 2024b; Fu et al., 2025b).
@@ 5.2 Long Context Benchmark
Long-context benchmark suites now serve as the standard yardstick for a broad range of evaluations—spanning RAG applications (Cheng et al., 2024; Qian et al., 2025; Fang et al., 2025), context-compression methods (Jiang et al., 2023; Pan et al., 2024; Jiang et al., 2024b), and many other long-sequence tasks. The development of long-context benchmarks like SCROLLS (Shaham et al., 2022) and LongBench (Bai et al., 2024b) has been crucial for evaluating LLMs’ extended sequence processing. However, these suites are also susceptible to data contamination, where inadvertent inclusion of evaluation samples in training data can inflate performance metrics (Xu et al., 2024d; Oren et al., 2023).
To address this, newer contamination-aware benchmarks like BAMBOO (Dong et al., 2024), LiveBench (White et al., 2024), and AcademicEval (Chen et al., 2025) have been introduced, often featuring continuously updated test sets to mitigate data leakage. A significant limitation, however, is their heavy reliance on laborious human effort for data collection and maintenance, rendering the process resource-intensive (Sainz et al., 2023b). Ensuring these updated benchmarks remain truly contamination-free is also challenging, as new data might still contain pre-existing knowledge (Yuan et al., 2024; Lee et al., 2024; Gao et al., 2025). Furthermore, undisclosed training data specifics for many LLMs make guaranteeing contamination absence in these benchmarks difficult (Modarressi et al., 2025; Qi et al., 2024).
@@ Appendix A Finetune Settings
We fine-tune four models on both datasets using one NVIDIA A100-SXM4-80GB GPU for 10 epochs. The finetuning process leverages the Unsloth library Daniel Han and team (2023) and incorporates LoRA Hu et al. (2021) with a rank of 32 and an alpha of 32. The initial learning rate is set at 2e-5. We adopt a per-device training batch size of 12 and set the gradient accumulation steps to 4 for LLaMA-3.1-8B and Qwen3-8B. Meanwhile, for Phi-4 and Qwen2.5-7B, we use a per-device training batch size of 8 and set the gradient accumulation steps to 6.

### v2

@@ Abstract
The increasing complexity of large language models (LLMs) raises concerns about their ability to “cheat” on standard Question Answering (QA) benchmarks by memorizing task-specific data. This undermines the validity of benchmark evaluations, as they no longer reflect genuine model capabilities but instead the effects of data leakage. While prior work has focused on detecting such leakage, little attention has been given to mitigating its impact and preserving the long-term utility of benchmarks. In this paper, we introduce LastingBench, a novel framework designed to continuously reinforce and safeguard existing benchmarks against knowledge leakage. LastingBench identifies leakage points in the context through perturbation, then rewrites the leakage points to counterfactual ones—disrupting memorization while preserving the benchmark’s original evaluative intent. Evaluations of state-of-the-art QA benchmarks show significant performance gaps, highlighting the efficacy of LastingBench in reducing memorization effects. LastingBench offers a practical and scalable solution to ensure benchmark robustness over time, promoting fairer and more interpretable evaluations of LLMs. Our code and data are available at https://github.com/Seriousss/LastingBench.
@@
Yixiong Fangthanks: Equal contribution.   Tianran Sun11footnotemark: 1   Yuling Shi   Min Wang   Xiaodong Guthanks: Correspondence: xiaodong.gu@sjtu.edu.cn Shanghai Jiao Tong UniversityUniversity of Pennsylvania {fangyixiong,seriousss,yuling.shi,xiaodong.gu}@sjtu.edu.cn minyun@seas.upenn.edu
@@ 1 Introduction
The rapid advancement of LLMs has introduced critical challenges to the reliability and validity of QA evaluation benchmarks Liu et al. (2024b); Wang et al. (2025a); Qian et al. (2025). Due to the opaque and large-scale nature of LLM training pipelines, these models often memorize parts of benchmark datasets Xu et al. (2024c); Balloccu et al. (2024); Geva et al. (2023); Deng et al. (2023); Cheng et al. (2025). This unintended data leakage enables models to “cheat” during evaluation, producing correct answers without genuine understanding. As a result, evaluation scores on standard benchmarks may no longer reflect genuine model capabilities, but instead represent latent memorization effects. This poses a serious threat to fair and meaningful measurement in NLP research Sainz et al. (2023a); Oren et al. (2023); Shi et al. (2023); Dekoninck et al. (2024b).
@@
In this paper, we propose LastingBench , a novel framework to reinforce and safeguard existing benchmarks against knowledge leakage, specifically targeting the long-context QA domain. Rather than continuously building new benchmarks, LastingBench enhances existing ones by identifying leakage points and rewriting them using counterfactuals, preserving the benchmark’s original purpose while disrupting model memorization. Our methodology begins with perturbation-based detection. By systematically perturbing the original context and the question, we assess whether a model relies on internal memorization rather than contextual clues. When leakage is detected, we locate critical evidence segments—the minimal context required to justify the answer—using enriched chain-of-thought (CoT) queries from an LLM. These segments are then rewritten into plausible yet semantically contradictory counterfactuals. This makes the benchmark robust against memorization while maintaining its reasoning challenge.
@@ 2.1 Study Design
1) Context Perturbation: We exclude the context and prompt the model to answer the original question based solely on the question. If the model can still provide the correct answer, i.e., $\mathcal{L}(q)=a^{\star}$, this suggests its reliance on internal knowledge, indicating knowledge leakage.
2) Question Perturbation: We either rephrase the original question $q$ into a semantically equivalent form $\tilde{q}$ or reformulate it into a logically contradictory version $q_{con}$, where the correct answer should conflict with the original. If the model answers the original question correctly but fails on the rephrased version, $\mathcal{L}(\tilde{q},\mathbf{C})\neq a^{\star}$, or provides the original answer even when the question is contradictory, $\mathcal{L}(q_{\text{con}},\mathbf{C})=a^{\star}$, this suggests dependence on memorized internal knowledge rather than contextual comprehension.
@@ 2.2 Experimental Setup
Studied Models We select a diverse set of models with varying parameter sizes and architectures for our experiments.(1) Both open-source models (Qwen2.5-32B-Instruct Qwen (2025), Llama-3.1-8B-Instruct et al (2024), Phi-4 Abdin et al. (2024)) and proprietary SOTA models including GPT-4o Hurst et al. (2024) and DeepSeek-v3 Liu et al. (2024a) are evaluated. (2) For the same model family, we also select models with different versions (e.g., Qwen3-32B vs. Qwen2.5-32B-Instruct) and parameter sizes(e.g., Qwen2.5-8B-Instruct vs. Qwen2.5-32B-Instruct), enabling multi-dimension comparisons.
@@ 2.3.2 Results by Question Perturbation
The results shown in Table 2 reveal varying levels of memorization across models. Notably, Claude-3-haiku shows the highest overlap on HotpotQA, with 40% of contradictory questions yielding the same answer as the original, despite the altered semantics. Similar trends are observed for models like Qwen2.5-32B-Instruct and Phi-4. These findings suggest that current LLMs heavily rely on memorized internal answers rather than reasoning over the given context.
@@ 2.4 Current Landscape and Trend
Our findings reveal that knowledge leakage is already widespread across long-context benchmarks, and the situation is deteriorating. Newer and larger models such as Qwen3 exhibit much higher leakage than their predecessors, as evidenced by drastic performance drops under question rephrasing. This trend suggests that mainstream benchmarks gradually no longer reliably reflect models’ genuine abilities—instead, they often measure memorization and result in inflated numbers. These findings highlight the urgent need for a pipeline to restore and defend benchmarks.
@@
[…]
@@ 3.2 Counterfactual Rewriting
To select the most effective counterfactual, we compute the conditional perplexity of the original question $q$ as:
@@
where $\text{PPL}(q)$ is the perplexity of the question $q$ answered without context, and $\text{PPL}(q\mid\mathbf{C}_{cf}^{(i)})$ measures perplexity given the i-th counterfactual. A higher CPPL indicates a stronger contradiction to the model’s internal knowledge. We select the counterfactual with the highest conditional perplexity:
@@ 4.1 Revised Dataset Details
We utilize DeepSeek-R1 as the leakage detecting model, evaluating with same methods in Section 2. Table 4 reports the percentages of revised (leaked) and unchanged entries for each dataset used in our experiments. Regarding the quality of the rewritten dataset, through manual inspection, we found that after rewriting, 99% of the contexts can fully support answering the question with the antifact answer. A genuine understanding of the full context should enable the model to generate the antifact answer. We acknowledge that a very small portion contained minor issues, such as unnecessary alterations to irrelevant parts. However, in most cases, the inaccurate modifications only involve distracting content that does not affect answering the question.
Dataset Revised Entries Unchanged Entries
@@ 4.2 Efficacy in Model Evaluation
As is displayed in Table 5, across all datasets, most models exhibit a consistent performance decline across different datasets. The performance drop is more pronounced on HotpotQA and 2WikiMQA, with HotpotQA showing the most significant decline (30% for Deepseek-v3). This aligns with our observations in §2.3.1 regarding context perturbation-based leakage detection. Among all evaluated models, Claude-3-Haiku (27% on HotpotQA) and the Qwen 3 series (30% on 2WikiMQA) show more pronounced performance degradation on the revised datasets, which is consistent with the leakage detection results presented in §2.3.2. This may be attributed to the fact that these models are relatively newer and therefore more likely to have been exposed to these datasets during the training stage, leading to potentially inflated performance. In addition, larger models tend to exhibit a higher degree of data memorization within their parameters. For example, GPT-4o and DeepSeek-v3 show a consistent decline on all the datasets. In contrast, our revised datasets provide a more trustworthy assessment of the models’ genuine long-context reasoning ability.
@@ 4.3 Comparative Analysis
Table 6 shows the model performance on HotpotQA. The results indicate that model performance on the randomly reformulated dataset falls short of the original, yet surpasses that on the counterfactual dataset, which suggests that our counterfactual rewriting method is more resistant to model cheating and poses a greater challenge, making it less likely to be memorized during pre-training. As a result, it offers a more faithful and enduring evaluation of the model’s long-context reasoning abilities.
@@ 4.4 Efficacy in Model Training
To examine whether our dataset defense effectively increases the learning complexity for language models and consequently reduces the likelihood of models memorizing information during the training stage, we conducted an empirical evaluation. Specifically, we fine-tuned four distinct models on both the original 2WikiMQA dataset and our revised dataset [^3] Details of our fine-tuning settings are provided in Section A. Introducing knowledge conflicts in datasets naturally creates greater complexity during the learning process. Ideally, this discourages models from memorizing knowledge.
As displayed in Figure 4, throughout the fine-tuning process, all models exhibit higher training loss on our revised dataset than on the original dataset, indicating that it exposes them to knowledge conflict situations. The consistently higher training loss values indicate that the revised data poses greater learning difficulty, which in turn mitigates the risk of future data leakage by making it less likely for models to memorize or retain such information.
4.5 Case Study
Table 7 illustrates how models may produce correct answers without genuinely reasoning over the provided context. In the first setting (Without Context), the model still predicts the correct year of appointment and the correct Olympic Games even though no supporting passages are given. This suggests that the model simply recalled memorized knowledge rather than performing contextual reasoning.
In contrast, the second part (Original/Rewritten Context) highlights the effect of counterfactual rewriting. When the original context explicitly contains the correct evidence (Marie of Hohenstaufen), the model outputs the right answer. However, after rewriting the critical evidence to state Joan of Arc, the model still produces the memorized answer, thus failing on the modified case. This discrepancy reveals that the earlier success was driven by leakage and memorization rather than true understanding of the context.
Without Context: After Russell D. Moore served at the Southern Baptist Theological Seminary, he became the President of The Ethics & Religious Liberty Commission (ERLC) in what year? GPT-4o: 2013 $\checkmark$
Professional cyclist Sara Symington competed in which Olympic Games held in Sydney, Australia? GPT-4o: 2000 $\checkmark$
Original/rewritten Context: Question: Who is the paternal grandmother of Marie of Brabant, Queen of France?
Original Context: … Marie of Brabant (13 May 1254 – 12 January 1322) was Queen of France from 1274 until 1285 as the second wife of King Philip III. Born in Leuven, Brabant, she was a daughter of Henry III, Duke of Brabant, and Adelaide of Burgundy. … Henry III of Brabant (c. 1230 – February 28, 1261, Leuven) was Duke of Brabant between 1248 and his death. He was the son of Henry II of Brabant and Marie of Hohenstaufen. Marie of Hohenstaufen
Rewritten Context: … Marie of Brabant (13 May 1254 – 12 January 1322) was Queen of France from 1274 until 1285 as the second wife of King Philip III. Born in Leuven, Brabant, she was a daughter of Henry III, Duke of Brabant, and Adelaide of Burgundy. … Henry III of Brabant (c. 1230 – February 28, 1261, Leuven) was Duke of Brabant between 1248 and his death. He was the son of Henry II of Brabant and Marie of Hohenstaufen. Joan of Arc Marie of Hohenstaufen Joan of Arc
Table 7: Case study showing knowledge leakage and counterfactual defense. Top: without any context, the model still gives correct answers (2013, 2000), revealing reliance on memorized knowledge. Bottom: in the Original Context, the correct answer (Marie of Hohenstaufen) is present, so the model answers correctly (✓), but this could be due to memorization rather than reasoning. In the Rewritten Context, we replace the key evidence with Joan of Arc. The model still outputs the original memorized answer, yielding an error (×), which confirms that the first success came from recall instead of contextual understanding.
@@ 5.1 Contamination Detection
The evaluation of Large Language Models (LLMs) is increasingly complicated by data contamination—the unintended overlap between training data and evaluation benchmarks (Fu et al., 2025; Cheng et al., 2025; Xu et al., 2024a). This issue can artificially inflate performance scores, misrepresenting LLMs’ true generalization capabilities. Vast pre-training datasets, often web-scraped, frequently include common benchmark samples, leading to deceptively high leaderboard scores that may not reflect real-world robustness (Tonmoy et al., 2024; Xu et al., 2024b; Fu et al., 2025).
@@ 5.2 Long Context Benchmark
Long-context benchmark suites now serve as the standard yardstick for a broad range of evaluations—spanning RAG applications (Cheng et al., 2024; Fang et al., 2025; Wang et al., 2025c), context-compression methods (Jiang et al., 2023; Pan et al., 2024; Jiang et al., 2024b), and many other long-sequence tasks (Wang et al., 2025b; Zhu et al., 2025; Liu et al., 2025b; Lan et al., 2025; Han et al., 2024; Liu et al., 2025a; Yang et al., 2025; Qiu et al., 2025; Zhang et al., 2025b, a). The development of long-context benchmarks like SCROLLS (Shaham et al., 2022) and LongBench (Bai et al., 2024b) has been crucial for evaluating LLMs’ extended sequence processing. However, these suites are also susceptible to data contamination, where inadvertent inclusion of evaluation samples in training data can inflate performance metrics (Xu et al., 2024d; Oren et al., 2023).
To address this, newer contamination-aware benchmarks have been developed, employing various strategies to combat data leakage. Some, like BAMBOO (Dong et al., 2024), LiveBench (White et al., 2024), and AcademicEval (Chen et al., 2025), feature continuously updated test sets. Others, such as VarBench, introduce dynamic variable perturbation to create more robust and generalizable evaluation scenarios (Qian et al., 2024). A significant limitation with manually updated benchmarks, however, is their heavy reliance on laborious human effort for data collection and maintenance, rendering the process resource-intensive (Sainz et al., 2023b). To mitigate this, approaches like Antileak-Bench aim to automatically construct benchmarks from updated real-world knowledge, offering a more scalable solution (Wu et al., 2024). Ensuring these updated benchmarks remain truly contamination-free is also challenging, as new data might still contain pre-existing knowledge (Yuan et al., 2024; Lee et al., 2024; Gao et al., 2025). Furthermore, undisclosed training data specifics for many LLMs make guaranteeing contamination absence in these benchmarks difficult (Modarressi et al., 2025; Qi et al., 2024).
@@ 6 Conclusion
Acknowledgment
This research is funded by the National Key Research and Development Program of China (Grant No. 2023YFB4503802) and the Natural Science Foundation of Shanghai (Grant No. 25ZR1401175).
@@ Appendix A Finetune Settings
We fine-tune four models on both datasets using one NVIDIA A100-SXM4-80GB GPU for 10 epochs. The fine-tuning process leverages the Unsloth library Daniel Han and team (2023) and incorporates LoRA Hu et al. (2021) with a rank of 32 and an alpha of 32. The initial learning rate is set at 2e-5. We adopt a per-device training batch size of 12 and set the gradient accumulation steps to 4 for LLaMA-3.1-8B and Qwen3-8B. Meanwhile, for Phi-4 and Qwen2.5-7B, we use a per-device training batch size of 8 and set the gradient accumulation steps to 6.